OpenStudy (anonymous):

Abby is touring China for four days. The car rental is $38.15 per day plus 12 1/2 cents per mile. If Abby drives 380.5 miles, how much does she owe the rental company?

OpenStudy (anonymous):

38.15 per day for 4 days = 38.15(4) = 152.60 12 1/2 cents per mile...same as 12.5 cents per mile.....for 380.5 miles = 12.5(380.5) = 4756.25 now add them 152.60 + 4756.25 = 4908.85 <== total cost